OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 10 and 12 and vertex at (11, -2). y = 2x^2 - 44x + 120 y = 2x^2 - 44x - 240 y = 2x^2 - 44x + 240 y = -2x^2 - 44x + 240

OpenStudy (anonymous):

I think it is c

OpenStudy (anonymous):

@jim_thompson5910 do you happen to know how to figure this out?

jimthompson5910 (jim_thompson5910):

is that a guess or do you have something to back you up?

OpenStudy (anonymous):

It is more of a guess

jimthompson5910 (jim_thompson5910):

it says "with zeros 10 and 12" so which equation has y = 0 when x = 10?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

How do I find that out?

jimthompson5910 (jim_thompson5910):

for each equation, plug in x = 10 and see which one results in y = 0

jimthompson5910 (jim_thompson5910):

choice A y = 2x^2 - 44x + 120 y = 2(10)^2 - 44(10) + 120 y = -120 so clearly x = 10 is NOT a root. So choice A is out

OpenStudy (anonymous):

Oh I see!